OpenStudy (anonymous):

what is the sp root of 50 plus the sp root of 32

OpenStudy (anonymous):

you mean: \[\sqrt{50}+\sqrt{32}\] ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it 1) 9 sq root of 2 2) sq root of 82 3) 6 4) 18

OpenStudy (anonymous):

=\[\sqrt{25*2}+\sqrt{16*2}\] =\[5\sqrt{2}+4\sqrt{2}\] =\[9\sqrt{2}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no problem